If F is Lipschitz function and g(x) is monotonically increasing on [a,b] then fog is bounded variation


1
Expert's answer
2022-03-07T17:18:01-0500

Let us consider a subdivision "a=x_0 \leq x_1 \leq ... \leq x_n=b" and the variation of "f\circ g" on this subdivision.

"V(f\circ g)=\sum_{i\leq n-1} |f\circ g(x_{i+1})-f\circ g(x_i)|"

First of all, as "f" is a Lipschitz, there is a constant "k" such that we have

"|f\circ g(x_{i+1})-f\circ g(x_i)| \leq k\cdot |g(x_{i+1})-g(x_i)|"

Now as "g" is increasing, we have "|g(x_{i+1})-g(x_i)|=g(x_{i+1})-g(x_i)"

From these two relations we deduce

"V(f\circ g)\leq k\cdot \sum|g(x_{i+1})-g(x_i)| =k\cdot \sum (g(x_{i+1})-g(x_i))"

The later sum is telescopic and gives "V(f\circ g)\leq k\cdot (g(b)-g(a))<+\infty"

Therefore, "f\circ g" is of bounded variation.